Heidelberg weekly meeting

03.410 (Physikalisches Institut, Heidelberg)


Physikalisches Institut, Heidelberg


We will have the meeting in HeiCONF: https://heiconf.uni-heidelberg.de/sch-nqu-dd2   Access Code: 858824

Vidyo will be used as a back-up. Room extension 10823320 or use this link.


Tentative minute taker list at https://docs.google.com/spreadsheets/d/15oo_a_1WPxPBOtlH0LDEKG1128y1O2lfswpUAvPCQOY/edit#gid=0

    • 14:00 14:45
      Round Table 45m

      Reports of work progress

      Speakers: Adrian Herkert (Physikalisches Institut, Uni Heidelberg), Alena Weber (Universität Heidelberg), Andre Schoening (University Heidelberg, Institute of Physiscs), Annie Meneses Gonzalez (PhD Student), Benjamin Weinläder (PI Heidelberg), David Maximilian Immig (Universität Heidelberg), Dohun Kim (Physikalisches Institut University of Heidelberg), Frank Meier Aeschbacher (Universität Heidelberg), Heiko Augustin (PI Uni Heidelberg), Lars Noehte (PI Heidelberg), Luigi Vigani (University of Oxford), Lukas Mandok (Physikalisches Institut Heidelberg), Marius Menzel, Sebastian Dittmeier (Physikalisches Institut, Universität Heidelberg), Sebastian Preuß (PI Uni Heidelberg), Thomas Theodor Rudzki (Physikalisches Institut Heidelberg)

      Minute Taker: David M. Immig

      Participants: Andre, David, Luigi, Thomas, Heiko, Alena, Alexander, Annie, Dohun, Florian, Jan, Lukas, Benny, Sebastian Preuß



      • -Presentation: 6-chip flexes plans: LTU in Lockdown up to January, redesign of new adapter board for readout
      • Molex Cable discussion: not twisted pair cables -> do our own? Cables designed for LVDS Signals up to 2 Gbit/s (concern of Andre it will not work->alternative for the data lines?
      • send 2nd probe card to PTSL


      • Magnet ramped to 2.7 T, the gradient in the magnetic field -> new measurements now how to operate at this field to get reasonable results
      • Milestone Magnet is operable is reached
      •  Former Monday DFG evaluation: referees seem to be happy, questions concerning Phase-II and thermal influence on pixel sensors ->asked for new Designer -> Alena
      • DFG funding 4+4 years, everything excepted (final approval in march)
      • TDR slightly delayed: fixing mistakes (found by yourself + referee comments)
      • Uppsala request: Bor coating on the backside of MP10 to measure neutrons, systematic study with americium (we do the measurement, they analyse)


      • a brief introduction to cross talk (presentation)
      • Reminder: Everyone should fill out Hackathon Doodle from Diddi


      • 2 Nights of Testbeam at HIT in HD with HITCounter Chip


      • FE-adapter board is taken over by Olivier (workshop): schematic done, layout on progress
      • Got access to KIP servers for simulations


      • working on simulations for paper


      • for every RUN2020 sensors type got one working sample (all sensors seems to be very noisy -> possible reason: VCasc2 connection from outside)


      • working on new test beam data sorter
      • working an abstract for BTTB9


      • working on new tiles with Florian
      • testing new MP10 (low breakdown voltage for 50um sensors: -20V to -25V)


      • worked with Dohun on new Tiles, designing new case for smaller tiles
      • David gave introduction to handle data


      • will order new bond wedges
      • ordering 6+6  more short HDIs with spacers
      • wafer probing card is not working -> try bond nicely looking sensor? take the risk to get (perhaps) working sensor on HDI before Xmas?





    • 14:45 14:55